# JEL Classification: C51, C81, D83, O13, Q12, Q54

### v1

@@ Appendix D Alternative EO Results
Note: Figure displays coefficients from event study regression (EquationΒ (4.1)). The dependent variable for Panel A is cumulative EVI, for Panel B is max EVI, for Panel C is mean EVI, and Panel D reproduces the results in the paper using median EVI values. The solid line shows coefficient estimates from the model, with the event year (the year immediately prior to the upazila having access to STRVs, indicated as -1) as the excluded category. Dotted lines represent confidence intervals calculated using standard errors clustered at the district-level.
Note: The figure presents a specification curve where each column represents a single difference-in-difference regression using the district panel remote sensing data but shifting the definition of treated (flood-prone). Numbers along the -axis are the number of districts in the treatment group. In the main paper, we define a district as flood-prone if a majority of the flood values in the district time series is above the median flood value in the entire data set , which results in 27 of 58 districts being treated. Here we vary the threshold for being considered treated. At the minimum, 14 districts fall above the median value in every year (left-most column) while at a maximum, 16 districts never rise above the median value in every year (right-most column). Significant and non-significant coefficients are designated at the top of the figure.
Cumulative | Max | Mean | Median | |
Panel A: Cumulative EVI | ||||
Post-STRV | 2035.3 | 2040.5 | 2035.3 | 2012.2 |
[1635.2,2435.5] | [1593.9,2487.2] | [1635.2,2435.5] | [1602.7,2421.7] | |
Flood Prone | -2752.2 | -3079.9 | -2752.2 | -2698.9 |
[-4601.7,-902.8] | [-4825.9,-1333.8] | [-4601.7,-902.8] | [-4510.1,-887.6] | |
Post Flood | 453.3 | 428.0 | 453.3 | 486.5 |
[-201.0,1107.6] | [-219.0,1074.9] | [-201.0,1107.6] | [-160.9,1133.9] | |
Observations | 1,280 | 1,280 | 1,280 | 1,280 |
Districts | 64 | 64 | 64 | 64 |
Adjusted R2 | 0.149 | 0.171 | 0.149 | 0.144 |
Panel B: Maximum EVI | ||||
Post-STRV | 0.0562 | 0.0531 | 0.0562 | 0.0560 |
[0.0449,0.0675] | [0.0418,0.0644] | [0.0449,0.0675] | [0.0444,0.0677] | |
Flood Prone | -0.102 | -0.0885 | -0.102 | -0.101 |
[-0.147,-0.0561] | [-0.134,-0.0428] | [-0.147,-0.0561] | [-0.145,-0.0559] | |
Post Flood | 0.00840 | 0.0146 | 0.00840 | 0.00854 |
[-0.0100,0.0268] | [-0.00340,0.0327] | [-0.0100,0.0268] | [-0.00974,0.0268] | |
Observations | 1,280 | 1,280 | 1,280 | 1,280 |
Districts | 64 | 64 | 64 | 64 |
Adjusted R2 | 0.225 | 0.171 | 0.225 | 0.221 |
Panel C: Mean EVI | ||||
Post-STRV | 0.0423 | 0.0423 | 0.0423 | 0.0418 |
[0.0340,0.0506] | [0.0331,0.0515] | [0.0340,0.0506] | [0.0333,0.0503] | |
Flood Prone | -0.0559 | -0.0631 | -0.0559 | -0.0549 |
[-0.0938,-0.0181] | [-0.0988,-0.0273] | [-0.0938,-0.0181] | [-0.0920,-0.0178] | |
Post Flood | 0.00916 | 0.00879 | 0.00916 | 0.00985 |
[-0.00443,0.0228] | [-0.00465,0.0222] | [-0.00443,0.0228] | [-0.00360,0.0233] | |
Observations | 1,280 | 1,280 | 1,280 | 1,280 |
Districts | 64 | 64 | 64 | 64 |
Adjusted R2 | 0.148 | 0.171 | 0.148 | 0.144 |
Note: The table displays regression results from a simple difference-in-difference specification in which we regression EVI on a post-treatment indicator equals 1 if the year is 2010 or later (post-STRV) and a treatment indicator equals 1 if the district is Flood Prone. We define a district as flood prone if a majority of the flood values in the district time series is above the median flood value in the entire data set . Column headers indicate the specific flood measure used in the regression. Standard errors are clustered at the district, with 95% confidence intervals displayed in brackets. |
Cumulative | Max | Mean | Median | |
Panel A: Cumulative EVI | ||||
STRV Availability | 0.615 | -0.264 | 0.615 | 0.516 |
[0.19,1.04] | [-1.60,1.07] | [0.19,1.04] | [0.30,0.73] | |
Flood | -1156 | -13065 | -26585 | -23070 |
[-1410,-901.1] | [-17811,-8317] | [-32443,-20727] | [-27135,-19005] | |
STRV Flood | -0.091 | 1.591 | -2.101 | -1.337 |
[-0.29,0.11] | [-2.55,5.73] | [-6.73,2.52] | [-4.46,1.79] | |
Observations | 1,280 | 1,280 | 1,280 | 1,280 |
Districts | 64 | 64 | 64 | 64 |
Within R2 | 0.753 | 0.697 | 0.753 | 0.750 |
Panel B: Maximum EVI | ||||
STRV Availability | -0.000 | -0.000 | -0.000 | 0.000 |
[-0.00,0.00] | [-0.00,0.00] | [-0.00,0.00] | [-0.00,0.00] | |
Flood | -0.0 | -0.3 | -0.6 | -0.6 |
[-0.04,-0.02] | [-0.43,-0.18] | [-0.81,-0.49] | [-0.70,-0.45] | |
STRV Flood | 0.000 | 0.000 | 0.000 | 0.000 |
[-0.00,0.00] | [-0.00,0.00] | [-0.00,0.00] | [-0.00,0.00] | |
Observations | 1,280 | 1,280 | 1,280 | 1,280 |
Districts | 64 | 64 | 64 | 64 |
Within R2 | 0.657 | 0.617 | 0.657 | 0.659 |
Panel C: Mean EVI | ||||
STRV Availability | 0.000 | -0.000 | 0.000 | 0.000 |
[0.00,0.00] | [-0.00,0.00] | [0.00,0.00] | [0.00,0.00] | |
Flood | -0.0 | -0.3 | -0.5 | -0.5 |
[-0.03,-0.02] | [-0.37,-0.17] | [-0.67,-0.42] | [-0.56,-0.39] | |
STRV Flood | -0.000 | 0.000 | -0.000 | -0.000 |
[-0.00,0.00] | [-0.00,0.00] | [-0.00,0.00] | [-0.00,0.00] | |
Observations | 1,280 | 1,280 | 1,280 | 1,280 |
Districts | 64 | 64 | 64 | 64 |
Within R2 | 0.750 | 0.696 | 0.750 | 0.746 |
Note: The table displays regression results from our two-way fixed effects (TWFE) specification with district and year fixed effects. The outcome variable is EVI and column headers indicate the specific flood measure used in the regression. Standard errors are clustered at the district, with 95% confidence intervals displayed in brackets. |
Note: Panel A shows the percentage of land area in the district dedicated to rice, normalized by the number of years before of after the introduction of STRV seed in that district. Panel B shows district-level values for EVI, normalized by the number of years before of after the introduction of STRV seed in that district. Year 0 is the year the seeds first became available in the district. We then plot non-parametric regressions to both the before and after data with confidence intervals.
Note: All panels show upazila-level values for the flood measure normalized by the number of years before of after the introduction of STRV seed in the district that the upazila is in. Year 0 is the year the seeds first became available in the district. We then plot non-parametric regressions to both the before and after data with confidence intervals.
Note: Figure displays coefficients from event study regression (EquationΒ (4.1)). The dependent variable is median EVI values in each upazila. The solid line shows coefficient estimates from the model, with the event year (the year immediately prior to the upazila having access to STRVs, indicated as -1) as the excluded category. Dotted lines represent confidence intervals calculated using standard errors clustered at the upazila-level.
Panel A: Difference-in-Difference | ||||
Cumulative | Max | Mean | Median | |
Post-STRV | 305.5 | 297.1 | 305.5 | 306.7 |
[280.3,330.6] | [269.1,325.2] | [280.3,330.6] | [282.6,330.8] | |
Flood Prone | -552.9 | -589.4 | -552.9 | -536.7 |
[-672.7,-433.1] | [-703.1,-475.8] | [-672.7,-433.1] | [-654.6,-418.8] | |
Post Flood | 26.0 | 40.9 | 26.0 | 23.2 |
[-19.9,71.9] | [-4.0,85.9] | [-19.9,71.9] | [-21.9,68.4] | |
Observations | 9,915 | 9,915 | 9,915 | 9,915 |
Upazilas | 503 | 503 | 503 | 503 |
Adjusted R2 | 0.152 | 0.164 | 0.152 | 0.147 |
Panel B: Two-Way Fixed Effects | ||||
Cumulative | Max | Mean | Median | |
STRV Availability | 0.122 | -0.064 | 0.122 | 0.103 |
[0.09,0.16] | [-0.16,0.03] | [0.09,0.16] | [0.08,0.12] | |
Flood | -180.6 | -2115.9 | -4153.1 | -3421.1 |
[-202.5,-158.6] | [-2509.0,-1722.8] | [-4657.9,-3648.2] | [-3739.8,-3102.3] | |
STRV Flood | -0.017 | 0.354 | -0.387 | -0.267 |
[-0.03,-0.00] | [0.09,0.62] | [-0.69,-0.08] | [-0.51,-0.02] | |
Observations | 9,915 | 9,915 | 9,915 | 9,915 |
Upazilas | 503 | 503 | 503 | 503 |
Within R2 | 0.587 | 0.504 | 0.587 | 0.571 |
Note: Panel A displays regression results from a simple difference-in-difference specification in which we regression EVI on a post-treatment indicator equals 1 if the year is 2010 or later (post-STRV) and a treatment indicator equals 1 if the upazila is Flood Prone. We define a upazila as flood prone if a majority of the flood values in the upazila time series is above the median flood value in the entire data set . Panel B displays regression results from our two-way fixed effects (TWFE) specification with upzaila and year fixed effects. Column headers indicate the specific flood measure used in the regression. Standard errors are clustered at the upazila, with 95% confidence intervals displayed in brackets. |
Appendix E Plot-Level Results
In the analysis relying on household data, we are concerned that our choice of the household, and not the plot, as the unit of analysis may be driving our results. If anything, conducting the analysis at the household-level is likely to create a downward bias, as we consider a household to have adopted if they grow STRVs on at least one plot, meaning many non-STRV plots cultivated by a household are considered βtreatedβ even though they are not. Because of this, our preference would be to conduct a plot-level analysis. However, we cannot track plots over time in the household panel, meaning we must treat the plot data as a repeated cross section.
We start by presenting summary statistics at the plot level which mirror the household level summary tables in the paper. Using plots as the unit of analysis, and controlling for household and year fixed effects, results are generally robust. In five of the eight regressions across both TWFE and TWFE-IV estimators, we reject that yields are equal for plots that experienced flooding and were planted with STRVs and plots that experienced flooding but were not planted with STRVs. These results should be interpreted with caution, as we cannot compare a plot with itself, meaning plot-level unobservables are left uncontrolled for.
@@ Appendix E Plot-Level Results
Rice Area (ha) | 0.642 | 0.719 | 0.032 | 0.651 |
(0.791) | (0.708) | (0.782) | ||
Cumulative Flooding | 2.393 | 1.944 | 0.000 | 2.343 |
(1.841) | (1.415) | (1.803) | ||
Max Flooding | 0.271 | 0.228 | 0.000 | 0.266 |
(0.159) | (0.131) | (0.157) | ||
Mean Flooding | 0.104 | 0.085 | 0.000 | 0.102 |
(0.080) | (0.062) | (0.078) | ||
Median Flooding | 0.085 | 0.070 | 0.000 | 0.084 |
(0.080) | (0.062) | (0.079) | ||
Village Flooding | 0.243 | 0.208 | 0.000 | 0.239 |
(0.123) | (0.104) | (0.121) | ||
Observations | 3598 | 457 | 4055 | |
Households | 1476 | 370 | 1499 | |
Note: The table displays summary statistics by adoption status and for the total sample with plot data. We also report the p-value on a t-test for the equality of means. |
2014 | 2017 | 2022 | 2014-2017 | 2017-2022 | 2014-2022 | ||
Yield (kg/ha) | 3786.629 | 3636.510 | 4172.232 | 0.009 | 0.000 | 0.000 | |
(1377.635) | (1636.447) | (1622.861) | |||||
Rice Area (ha) | 0.701 | 0.670 | 0.569 | 0.297 | 0.000 | 0.000 | |
(0.944) | (0.622) | (0.738) | |||||
Cumulative Flooding | 2.438 | 2.297 | 2.286 | 0.036 | 0.867 | 0.033 | |
(1.803) | (1.778) | (1.831) | |||||
Max Flooding | 0.260 | 0.251 | 0.290 | 0.097 | 0.000 | 0.000 | |
(0.154) | (0.161) | (0.152) | |||||
Mean Flooding | 0.106 | 0.100 | 0.099 | 0.036 | 0.867 | 0.033 | |
(0.078) | (0.077) | (0.080) | |||||
Median Flooding | 0.087 | 0.082 | 0.082 | 0.062 | 0.835 | 0.123 | |
(0.079) | (0.078) | (0.080) | |||||
Village Flooding | 0.234 | 0.225 | 0.262 | 0.054 | 0.000 | 0.000 | |
(0.120) | (0.127) | (0.114) | |||||
Observations | 1412 | 1446 | 1197 | ||||
Households | 1412 | 1446 | 1197 | ||||
Note: The table displays summary statistics by year. We also report the p-value on a t-test for the equality of means. |
Panel A: Two-Way Fixed Effects | |||||
Cumulative | Max | Mean | Median | Village | |
STRV Adoption | 0.097 | 0.026 | 0.097 | 0.093 | -0.054 |
[-0.10,0.30] | [-0.26,0.32] | [-0.10,0.30] | [-0.09,0.27] | [-0.40,0.29] | |
Flood | -0.432 | -3.754 | -9.926 | -5.086 | -5.383 |
[-0.57,-0.29] | [-4.83,-2.68] | [-13.19,-6.66] | [-7.34,-2.83] | [-6.83,-3.93] | |
STRV Flood | -0.013 | 0.124 | -0.295 | -0.173 | 0.538 |
[-0.09,0.07] | [-1.09,1.34] | [-2.11,1.52] | [-1.53,1.19] | [-0.86,1.93] | |
Observations | 3,865 | 3,865 | 3,865 | 3,865 | 3,865 |
Households | 1,488 | 1,488 | 1,488 | 1,488 | 1,488 |
Within R2 | 0.183 | 0.195 | 0.183 | 0.162 | 0.202 |
Panel B: Two-Way Fixed Effects - IV | |||||
Cumulative | Max | Mean | Median | Village | |
STRV Adoption | -0.254 | -1.533 | -0.254 | 1.469 | -1.332 |
[-0.97,0.46] | [-2.72,-0.35] | [-0.97,0.46] | [0.64,2.30] | [-2.23,-0.44] | |
Flood | -0.374 | -3.582 | -8.604 | -3.439 | -4.432 |
[-0.52,-0.23] | [-4.95,-2.22] | [-11.90,-5.31] | [-6.09,-0.79] | [-5.97,-2.89] | |
STRV Flood | 0.010 | 5.796 | 0.224 | -7.517 | 6.053 |
[-0.21,0.23] | [1.41,10.18] | [-4.76,5.21] | [-12.77,-2.27] | [2.64,9.47] | |
Observations | 3,865 | 3,865 | 3,865 | 3,865 | 3,865 |
Households | 1,488 | 1,488 | 1,488 | 1,488 | 1,488 |
Within R2 | 0.058 | 0.034 | 0.058 | 0.006 | 0.049 |
Note: Panel A displays regression results from our two-way fixed effects (TWFE) specification with household and year fixed effects but data at the plot-level. Panel B displays regression results from our TWFE model in which we instrument for adoption using the householdβs past history of flooding. Column headers indicate the specific flood measure used in the regression. Standard errors are clustered at the household, with 95% confidence intervals displayed in brackets. |

### v2

@@ E.2 Upazila-Level Results
Note: Figure displays coefficients from district-level event study regressions that excludes 19 coastal districts. The dependent variable for Panel A is cumulative EVI, for Panel B is max EVI, for Panel C is mean EVI, and Panel D reproduces the results in the paper using median EVI values. The solid line shows coefficient estimates from the model, with the event year (the year immediately prior to the upazila having access to STRVs, indicated as -1) as the excluded category. Dotted lines represent confidence intervals calculated using standard errors clustered at the district-level.
Note: The figure presents results from district-level TWFE regressions that excludes 19 coastal districts. Each panel in the figure displays coefficient estimates and confidence intervals on the interaction of STRVs and a flood metric for a specific EVI measure (cumulative, max, mean, median). Specifications are sorted based on coefficient size from smallest (left) to largest (right). The gray diamonds below the coefficients indicate which combination of quantile and flood window was used in the regression.
Note: The figure presents results from district-level TWFE regressions that excludes 19 coastal districts. Each panel in the figure displays coefficient estimates and confidence intervals on the interaction of STRVs and a flood metric for a specific EVI measure (cumulative, max, mean, median). Specifications are sorted based on coefficient size from smallest (left) to largest (right). The gray diamonds below the coefficients indicate which combination of quantile and flood window was used in the regression.
Note: The figure presents results from district-level TWFE regressions that excludes 19 coastal districts. Each panel in the figure displays coefficient estimates and confidence intervals on the interaction of STRVs and a flood metric for a specific EVI measure (cumulative, max, mean, median). Specifications are sorted based on coefficient size from smallest (left) to largest (right). The gray diamonds below the coefficients indicate which combination of quantile and flood window was used in the regression.
Note: Figure displays coefficients from upazila-level event study regressions. The dependent variable for Panel A is cumulative EVI, for Panel B is max EVI, for Panel C is mean EVI, and Panel D reproduces the results in the paper using median EVI values. The solid line shows coefficient estimates from the model, with the event year (the year immediately prior to the upazila having access to STRVs, indicated as -1) as the excluded category. Dotted lines represent confidence intervals calculated using standard errors clustered at the district-level.
Note: The figure presents results from upazila-level TWFE regressions. Each panel in the figure displays coefficient estimates and confidence intervals on the interaction of STRVs and a flood metric for a specific EVI measure (cumulative, max, mean, median). Specifications are sorted based on coefficient size from smallest (left) to largest (right). The gray diamonds below the coefficients indicate which combination of quantile and flood window was used in the regression.
Note: The figure presents results from upazila-level TWFE regressions. Each panel in the figure displays coefficient estimates and confidence intervals on the interaction of STRVs and a flood metric for a specific EVI measure (cumulative, max, mean, median). Specifications are sorted based on coefficient size from smallest (left) to largest (right). The gray diamonds below the coefficients indicate which combination of quantile and flood window was used in the regression.
Note: The figure presents results from upazila-level TWFE regressions. Each panel in the figure displays coefficient estimates and confidence intervals on the interaction of STRVs and a flood metric for a specific EVI measure (cumulative, max, mean, median). Specifications are sorted based on coefficient size from smallest (left) to largest (right). The gray diamonds below the coefficients indicate which combination of quantile and flood window was used in the regression.
Appendix F Household-Level Robustness Checks
A third potential concern regarding our results may be driven by our use of EVI as a proxy for yields and seed dissemination data as a proxy for adoption. Rice is not cultivated at the district-level but by individual households on small fields. District-level EO measures may be noisy or biased proxies of the on-the-ground household reality. To demonstrate robustness of our district-level results, we estimate TWFE regressions using household data on yields and adoption and EO data on floods experienced by that households. In this appendix we discuss the value and limitations of the three-year Rice Monitoring Survey (RMS) panel as well as summary statistics from the data and finally results of robustness checks.
F.1 Household Survey Data
The RMS was a Gates Foundation-funded project designed to capture varietal turnovers over time in Bangladesh, India, and Nepal. The data was originally collected as two waves of a panel in 2014 and 2017 (Yamano, 2017). Households were selected following a clustered random sampling procedure to ensure the overall survey was representative of rice growing regions of each country. The plan was to collect a third wave in 2020, but data collection efforts were delayed due to COVID until 2022 and only Bangladesh was revisited. In total, 1,500 households were part of the initial sampling frame in Bangladesh. The RMS was able to follow-up with 1,490 households in 2017 and 1,484 households in 2022. This gives us an attrition rate of over eight years.
We construct an unbalanced panel of households that cultivated rice in Aman season. This gives us a total of 3,865 observations from 1,488 households. Our primary outcome of interest is rice yield at the household level, which we construct by summing up rice harvest (kg) on all plots and dividing by the sum of area (ha) for all rice plot. We consider a household as having adopted STRVs if they report planting any flood tolerant rice variety on at least one plot. To measure the incidence of flooding faced by the household, we combine household GPS location with the EO flood maps we generate. While we would ideally use plot locations to match with flood data, only household GPS locations were collected in 2014 and 2017. Thus, we only know the precise location of plots in 2022.
TableΒ F1 provides summary statistics for key variables across the three years of the panel. The first three columns of the table report means and standard deviations for key variables in each year. The final three columns of the table report -values on year-to-year comparisons of means using a -test that allows for unequal variance. Statistically significant differences exist for almost every variable in almost every year-to-year comparison. Yields in 2014 and 2017 were around tons per hectare and increased to tons per hectare in 2022. Adoption was flat at eight percent of households for the first two years in the panel (in 2014, households adopted STRVs, in 2017, households adopted). Adoption grew substantially in the intervening five years, with ( households) adopting STRVs in 2022. Rice area has steadily declined over time and while statistically significant, the magnitude of the decline is not very large. In 2014, households had hectares in rice, which fell by a tenth of a hectare to in 2022.
TableΒ F2 presents summary statistics by adoption status and for the dataset overall. Households that adopt STRVs report higher yields in the years that they adopt compared to households not growing STRVs in a given year. Households can and do dis-adopt STRVs, though given how low adoption rates were in 2014 and 2017, dis-adoption is much less common that adoption. Recall also that we classify a household as adopting if they use STRVs on at least one of their rice plot. Thus, yields for adopters includes yields from both STRV and non-STRV seeds. The difference in yields between adopters and non-adopters suggests that the adoption decision is not random and may be correlated with household or farm unobservables, such as farmer ability or quality and location of plots. That said, we do not see significant differences in terms of rice area between adopters and not adopters.
F.2 Household Empirical Method
In the case of the household data, a event study or DID does not make sense because treatment (adoption) can turn on and off again as households choose to adopt or dis-adopt. So we must limit our analysis to a TWFE regression similar to our district-level framework. We estimate the effects of STRV adoption by a household on rice yields as:
where indexes households and indexes time. The two terms and denote household and time fixed effects, meaning that only within-household time variation in relative yields remains. The household fixed effects control for all household-specific time-invariant variation and the time fixed effects control for all time-specific household-invariant variation.
In terms of identification, modelΒ (F.1), like modelΒ (4.3) relies on household and time fixed effects to identify changes in a householdβs yields resulting from a change in the householdβs adoption status. The model is identified only if the adoption decision is not drive by household-specific, time-varying events. We believe that adoption is plausibly exogenous and primarily driven either by time-invariant household characteristics like risk preferences and farming skill or by availability of seed, which changes over time but not uniquely for a household. In all models we cluster standard errors at the level of the unit of observation.
F.3 Econometric Evidence
FigureΒ F1 reports results using a binary indicator for if flooding fell within the window (panel A) and using the number of days in the flood window (panel B). Unlike results from our disaggregation to upazila, the household-level results are similar to the district-level TWFE results. Again, the vast majority of regressions are not significant or negative and significant. In both panels, slightly more than of regressions produce positive and significant results. The flood metrics that matter for the household, though, are from lower quantiles than at the district level, meaning that when we use the pixel-level fractional flood index (as opposed to district-level averages of the pixel-level data) a greater degree of inundation above baseflow is required in order to generate a Goldilocks flood.
In terms of checking the robustness of our main results, the preponderance of evidence from the household-level analysis supports our conclusions from the district-level EO analysis. STRVs have a positive and significant effect on yields when flooding is βjust right.β At a practical level, for those looking to address the Goldilocks problem for other agricultural technologies, our household- v.s. district-level findings suggest that what qualifies as a βjust rightβ stress event will vary based on the unit of analysis and the aggregation (or disaggregation) of the data.
2014 | 2017 | 2022 | 2014-2017 | 2017-2022 | 2014-2022 | ||
Yield (kg/ha) | 3787 | 3637 | 4172 | 0.009 | 0.000 | 0.000 | |
(1378) | (1636) | (1623) | |||||
STRV Adoption | 0.081 | 0.081 | 0.213 | 0.997 | 0.000 | 0.000 | |
(0.272) | (0.272) | (0.409) | |||||
Rice Area (ha) | 0.720 | 0.697 | 0.619 | 0.468 | 0.005 | 0.003 | |
(0.950) | (0.619) | (0.750) | |||||
Observations | 1376 | 1389 | 1100 | ||||
Households | 1376 | 1389 | 1100 | ||||
Note: The table displays means for key variables by panel year. Standard deviations are in parentheses. We also report -values on a -test for the equality of means between adopters and non-adopters. |
@@ F.3 Econometric Evidence
Observations | 3408 | 457 | 3865 | |
Households | 1460 | 370 | 1488 | |
Note: The table displays means for key variables by adoption status and for the total sample. Standard deviations are in parentheses. We also report -values on a -test for the equality of means between adopters and non-adopters. |
Note: The figure presents results from household-level TWFE regressions. The figure displays coefficient estimates and confidence intervals on the interaction of household adoption of STRVs and a flood metric on log yield. The top panel reports on results using a binary if flooding was in the window while the bottom panel reports on results using the number of days in the window. Specifications are sorted based on coefficient size from smallest (left) to largest (right). The gray diamonds below the coefficients indicate which combination of quantile and flood window was used in the regression.